OpenStudy (kaylardz0405):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 5 is rolled? A. 333 B. 500 C. 167 D. 667

OpenStudy (kaylardz0405):

@dtan5457 @jagr2713

OpenStudy (kaylardz0405):

@limecake08

OpenStudy (kaylardz0405):

will fan and medal

OpenStudy (anonymous):

It would be A

OpenStudy (kaylardz0405):

why

OpenStudy (anonymous):

numbers greater than 4, 5 and 6 chance per roll of getting 5 = 1/6, of getting 6 = 1/6 of 5 or 6 = 1/6 + 1/6 = 1/3 1000 times = 1000*1/3 = 333

OpenStudy (kaylardz0405):

ok thank you

OpenStudy (anonymous):

No problem!